Write an equation for the parabola whose vertex is at (-8,4) and passes through (-6,-2).
Thanks

Answer by venugopalramana(3286)

Write an equation for the parabola whose vertex is at (-8,4) and passes through (-6,-2).
STANDARD EQN. IS
(Y-K)^2=4A(X-H)..WHERE
VERTEX IS (H,K)....SO H=-8...K=4
(Y-4)^2=4A(X+8)...IT GOES THROUGH (-6,-2).SO.
36=4A*2
A=36/8=4.5
(Y-4)^2=18(X+8)